ok ive tried to figure this out over and over and i just dont get it! how do u put a problem solved for slope (y=mx+b) into so called standard form?? it makes no sence!
Answer by longjonsilver(2297)

a straight line equation standard form is the y=mx+b format.

So y=-2x+5 is standard form whereas 3y = 6x-12 is not standard form --> y = 2x-4 is though.

Jon.